# Sylva: Tailoring Personalized Adversarial Defense in Pre-trained Models via Collaborative Fine-tuning

### v1

@@ 3.3. Cooperative Game-based Fine-tuning
After Phase 1, each client fuses the global LoRA with the pre-trained model to obtain a robustly generalized backbone. Adding a personalized classifier enables effective adversarial defense. However, we believe that benign accuracy can be further improved without sacrificing robustness.
3.3.1. Shapley Game for Layers
@@
In Phase 2, each client independently trains using its own data in parallel, further fine-tuning its personalized model. Our goal is to select $p$ layers from the $L$ layers of the robust model for training on benign data, while freezing the remaining layers, to maximize accuracy improvements and minimize robustness degradation. This problem is framed as a cooperative game, with Shapley values (Shapley, 1953) used to quantify the marginal contribution of each layer and optimize the overall value function.
Let $\mathbf{L}$ represent the set of model layers, and $\mathbf{S}$ denote one subset of layers selected for training while freezing the rest. We define two sensitivity losses as
(16) $\mathcal{L}_{rob}(\mathbf{S})=\mathcal{L}_{CE}(\mathcal{F}(w_{i}^{\{\mathbf{S}\}},x^{adv}),y)-\mathcal{L}_{CE}(\mathcal{F}(w_{i},x^{adv}),y),$
(17) $\mathcal{L}_{acc}(\mathbf{S})=\mathcal{L}_{CE}(\mathcal{F}(w_{i}^{\{\mathbf{S}\}},x),y)-\mathcal{L}_{CE}(\mathcal{F}(w_{i},x),y),$
to quantify the impact of training $\mathbf{S}$: $\mathcal{L}_{rob}$ measures changes in loss for adversarial samples, and $\mathcal{L}_{acc}$ measures changes in loss for benign samples. A smaller $\mathcal{L}_{rob}$ indicates that the selected layers are less sensitive to robustness, meaning fine-tuning them minimally affects robustness. Conversely, a larger $\mathcal{L}_{acc}$ suggests that the selected layers have potential for further optimization on benign data. Based on these metrics, we define the value function in the cooperative game as
(18) $v(\mathbf{S})=\mathcal{L}_{acc}(\mathbf{S})-\beta\cdot\mathcal{L}_{rob}(\mathbf{S}),$
where $\beta$ is a hyperparameter used to adjust the value in the cooperative game. Subsequently, we can calculate the Shapley value for each layer $l$ as its average marginal contribution, denoted as
(19) $\phi_{l}=\sum_{\mathbf{S}\subseteq\mathbf{L}\setminus\{l\}}\frac{|\mathbf{S}|!(L-|\mathbf{S}|-1)!}{L!}\left[v(\mathbf{S}\cup\{l\})-v(\mathbf{S})\right].$
This parameter accounts for each participating layer’s marginal contribution across all possible combinations of participants and computes the average. Finally, we select $p$ layers, denoted as $\mathbf{P}=\{l_{1},l_{2},\ldots,l_{p}\}$, where $\phi_{l_{1}}\geq\phi_{l_{2}}\geq\cdots\geq\phi_{l_{p}}$. By freezing the other layers and fine-tuning only these selected layers using benign data, the model’s accuracy can be effectively improved.
3.3.2. Monte Carlo Sampling Optimization
While the cooperative game approach leveraging Shapley values effectively identifies the optimal layers for fine-tuning to enhance the model’s benign accuracy, its computational complexity is prohibitively high. Specifically, for each layer, it requires evaluating all subsets of the remaining layers and computing the value function, resulting in a complexity of $O(2^{L}\times L)$. To address this, we plan to adopt Monte Carlo sampling to optimize computational cost.
We randomly generate $B$ permutations of the model layers, denoted as $\pi^{(1)},\pi^{(2)},\ldots,\pi^{(B)}$. For each permutation $\pi^{(b)}$, we identify the position of layer $l$ and define the set of all layers preceding $l$ in that permutation as $S_{l}^{(b)}$. The marginal contribution of layer $l$ when added to $S_{l}^{(b)}$ is then represented as
(20) $\phi_{l}\approx\frac{1}{B}\sum_{b=1}^{B}\left[v(\mathbf{S}_{l}^{(b)}\cup\{l\})-v(\mathbf{S}_{l}^{(b)})\right].$
As the number of samples $B$ increases, the calculated contribution values converge to the results of the Shapley value game. The computational complexity is reduced to $O(B\times L)$, with the error decreasing at a rate of $O(1/\sqrt{B})$ as $B$ grows.
3.4. The Workflow of Sylva
Algorithm 1 Sylva’s Adversarial Training Process
@@ 3.4. The Workflow of Sylva
1:  Input: Local training epoch $T_{1}$, aggregation epoch $T_{2}$, local benign data $\mathcal{D}_{i}$, global model $w_{g}(w_{g}^{L},w_{g}^{P},w_{g}^{C})$, and local model $w_{i}(w_{i}^{L},w_{i}^{P},w_{i}^{C})$, where $i\in\{1,2,\dots,N\}$.
@@
3:  In Phase 1:
4:  Initialize: Server’s model $w_{g}$ and clients’ models $w_{i}$.
5:  for $1$ to $T_{2}$  do
6:     for client $i\in\{1,2,\dots,N\}$ parallel do
7:        Client $i$ downloads the global LoRA model as $w_{i}^{L}$. Combine $w_{i}^{L}$, $w_{i}^{P}$ and $w_{i}^{C}$ to form $w_{i}$;
8:        for 1 to $T_{1}$ do
9:            for $x\in\mathcal{D}_{i}$ do
10:               $x_{adv}=\text{PGD}(w_{i},x)$;
11:               Train and calculate loss $\mathcal{L}$ with Eq. 11;
12:               $w_{i}^{L},w_{i}^{C}\leftarrow\text{SGD}(w_{i},\mathcal{L})$;
@@
14:        end for
15:        Each client upload LoRA model $w_{i}^{L}$ to server;
16:     end for
17:     Server aggregate LoRA model with Eq. 15;
18:  end for
19:  In Phase 2:
20:  for client $i\in\{1,2,\dots,N\}$ parallel do
21:     Fuse $w_{i}^{L}$, $w_{i}^{P}$, and $w_{i}^{C}$ to form the model $w_{i}$;
22:     for layer $l\in w_{i}$ do
23:         Calculate marginal contribution with Eq. 20;
24:     end for
25:     Select layers $\mathbf{P}=\{l_{1},l_{2},\dots,l_{p}\}$, $\phi_{l_{1}}\geq\phi_{l_{2}}\geq\cdots\geq\phi_{l_{p}}$;
26:     Freeze layers except $\mathbf{P}$ and train model $w_{i}$ by $\mathcal{D}_{i}$;
27:  end for
Algorithm 1 outlines the Sylva defense algorithm. Each client $i$ holds its private benign dataset $\mathcal{D}_{i}$. A global model $w_{g}$ and a local model $w_{i}$, both utilizing LoRA for fine-tuning, are constructed and deployed.
In Phase 1, personalized adversarial fine-tuning is implemented with the following steps:
  • Initialize the server model $w_{g}$, and allow clients to download the LoRA model’s parameters $w_{g}^{L}$ from the server. Each client then constructs its local model $w_{i}$. (Line 4-7)
  • Each client generates adversarial examples from its local dataset $\mathcal{D}_{i}$ using the PGD method. Then, the client trains locally for $T_{1}$ epochs using Eq. 11, only updating its LoRA parameters $w_{i}^{L}$ and classifier parameters $w_{i}^{C}$. (Line 8-14)
  • Clients upload their updated $w_{i}^{L}$ parameters to the server, which aggregates them using Eq. 15 to update the global LoRA parameters $w_{g}^{L}$. This process is repeated for $T_{2}$ rounds. (Line 15-18)
After completing Phase 1, each client possesses the shared global LoRA model $w_{i}^{L}$, the frozen pre-trained backbone $w_{i}^{P}$, and a personalized classifier $w_{i}^{C}$. Phase 2 involves cooperative game-based fine-tuning, which proceeds as follows:
  • Each client fuses the updated $w_{i}^{L}$, $w_{i}^{P}$, and $w_{i}^{C}$ from Phase 1 to construct the complete local model $w_{i}$. It then calculates the marginal contribution $\phi_{l}$ for each layer $l$ using Eq. 20, quantifying the relative contribution of each layer. (Line 21-24)
  • Based on the computed marginal contributions, the top-$p$ layers with the highest contribution values are selected to form the set $\mathbf{P}$. (Line 25)
  • Using the client’s local benign data, all layers outside the set $\mathbf{P}$ are frozen, while the layers within $\mathbf{P}$ are fine-tuned to enhance the model’s benign accuracy. (Line 26)
Through this workflow, each client obtains a personalized large model tailored to its data distribution, achieving high accuracy on benign data while effectively resisting adversarial attacks.
Table 2. Performance comparison of Sylva and baseline under different datasets and attack algorithms (ViT-B/16)
@@
Sylva(Ours) 26.95 27.63 26.23 28.95 28.70
4. Experiments
Throughout our experimental evaluation, we explore the performance of Sylva defense under diverse distributed environments and attack scenarios. We compare Sylva with baselines to highlight its personalized defense advantages and assess its feasibility and adaptability across different settings, addressing four key research questions.
  • RQ1: How does Sylva perform against different attack algorithms in terms of robustness and accuracy?
  • RQ2: How does Sylva handle varying heterogeneous data distributions across diverse clients?
  • RQ3: How does Sylva achieve efficiency in communication and computation in distributed environments?
  • RQ4: How do Sylva’s hyperparameters influence its overall robustness and accuracy performance?
4.1. Experimental Setup
@@
Datasets and models. This study leverages datasets widely recognized in adversarial training, including CIFAR-10 (Krizhevsky et al., 2009), STL-10 (Coates et al., 2011), and GTSRB (Stallkamp et al., 2011), as introduced in Section 2. To better align with tasks involving large models, the more complex CIFAR-100 dataset is also incorporated. For model selection, we utilize large vision models based on transformers, specifically ViTs (Dosovitskiy et al., 2021) pre-trained on the ImageNet dataset (Russakovsky et al., 2015). These models are categorized into three scales—ViT-T/16, ViT-B/16, and ViT-L/16—based on the number of transformer layers.
@@
Figure 7. Comparison of adversarial training performance under different heterogeneity levels (ViT-B/16)
@@
Attacks. We simulate both white-box and gray-box attacks using recent and widely adopted image attack algorithms. For white-box attacks, we employ the classic FGSM algorithm (Goodfellow et al., 2014) and its variants, alongside the more robust PGD algorithm (Madry et al., 2018). To explore subtle perturbations, we include the SparseFool attack (Modas et al., 2019), which modifies only a few pixels. For gray-box attacks, we utilize the PAP algorithm (Ban and Dong, 2022), which generates universal perturbations tailored for pre-trained models and remains effective against fine-tuned models without requiring knowledge of downstream tasks.
@@
Baselines. We use several standard defense algorithms for adversarial training: PGD-AT (Madry et al., 2018), which employs the PGD algorithm to generate adversarial samples and enhance model robustness; TRADES (Zhang et al., 2019), which optimizes the loss to balance accuracy on benign data and robustness against adversarial attacks; and Gen-AF (Zhou et al., 2024), a fine-tuning approach for adversarial training on pre-trained models, maintaining robustness while improving accuracy. In the distributed scenario, we employ FedAvg (McMahan et al., 2017) and FedProx (Li et al., 2020b), with FedProx addressing client data heterogeneity. We integrate them with various defense algorithms, establishing them as baseline approaches for comparison. Additionally, we incorporate DBFAT (Zhang et al., 2023b), designed specifically for non-IID settings, into our baseline approaches. We also include two personalized federated learning algorithms for comparison. (Allouah et al., 2024) proposes addressing Byzantine attacks in distributed training by customizing the loss function for each client, which we refer to as Per-Adv. (Wagner et al., 2024) introduces a method that computes LoRA-based trust weights among clients during large model fine-tuning to achieve personalized updates, which we denote as Per-LoRA. It is worth mentioning that we incorporate the TRADES loss into both personalized algorithms, thereby improving their alignment with the threat model and experimental tasks defined in this study.
@@
Implementation. For data heterogeneity, client data is partitioned using a Dirichlet distribution (Li et al., 2020a) with 15 clients and a Dirichlet parameter of 10. In Phase 1, $\gamma$ and $\epsilon$ are set to 0.9, $\lambda_{1}$ to 20, $\lambda_{2}$ to 0.001, and $k$, the number of clients for weight aggregation, is set to 5. In Phase 2, $\beta$ is set to 5, $B$ to 300, and $p$, the proportion of layers selected for training, to 3% of the total layers. The learning rate is 0.005. Federated adversarial training uses 5 local epochs, with ViT-T performing 30 cloud aggregation rounds, while ViT-B and ViT-L performing 50. Experiments are conducted on 4 NVIDIA A100 GPUs using the Ray framework (Moritz et al., 2018) for multi-process client simulation. To assess Sylva’s efficiency in real-world scenarios, we conduct experiments on five real devices, which are used to simulate edge devices in various practical applications. These include three consumer-grade GPUs: GeForce RTX 4090, 3090, and 2080Ti. Additionally, we test two edge computing devices for autonomous driving. One is the GeForce RTX 3060 in the Apollo D-KIT Advanced (Baidu, 2022), a kit for autonomous driving simulation, and the other is the Jetson AGX Orin (NVIDIA, 2021), which is commonly used in real-world vehicles (NVIDIA, 2024b, a).
4.2. Performance of Adversarial Training (RQ1)
@@
We evaluate on four datasets, reporting average performance across all clients (Table 2). In heterogeneous settings, non-IID-aware baselines like FedProx variants and DBFAT show relatively better performance. Sylva consistently outperforms across diverse datasets, excelling on both benign and adversarial samples. On datasets like CIFAR-10, it shows substantial gains over both traditional federated adversarial training and personalized methods. Specifically, Sylva leads to an accuracy increase of up to 50.4% and a robustness enhancement of 29.5%, demonstrating its superior capability in enhancing both performance and resilience against adversarial attacks. When compared to the baselines of the two personalized algorithms, Sylva continues to deliver significant improvements, achieving an average increase of 48.3% in benign accuracy and 17.5% in robustness. This performance boost can be attributed to Sylva’s greater adaptability, as it effectively balances generalization across tasks with maximizing personalization for individual clients, thereby enhancing both accuracy and robustness. While the improvements on more complex datasets like CIFAR-100 are comparatively smaller, they remain significant, highlighting Sylva’s adaptability and effectiveness in varied scenarios. Furthermore, we evaluate Sylva on models of varying scales, with detailed results provided in Appendix B. Sylva consistently delivers strong performance across all model scales.
@@
Answer to RQ1: By leveraging its personalized design, Sylva enables each client’s model to effectively adapt to its specific data distribution, achieving superior accuracy and robustness even in highly heterogeneous environments.
4.3. Performance of Different Heterogeneous Distributions (RQ2)

### v3

@@ III-D Forward-gating Selection Fine-tuning
After Phase 1, each client fuses the global LoRA-FA with the pre-trained model to obtain a robustly generalized backbone. Adding a personalized classifier enables effective adversarial defense. However, we believe that benign accuracy can be further improved without sacrificing robustness.
@@
In Phase 2, each client independently trains using its own data in parallel, further fine-tuning its personalized model. Our goal is to select $p$ layers from the $L$ layers of the robust model for training on benign data, while freezing the remaining layers, to maximize accuracy improvements and minimize robustness degradation. Previous studies jointly compute the robustness and accuracy losses for each layer to determine the corresponding layer selection [28, 48], but this method involves high computational complexity.
As shown in fig. 6, for each client $i$, we first fuse the locally trained weights $w_{i}^{L}$ obtained from Phase 1 with its backbone model. After attaching a personalized classifier, we obtain the fine-tuned model $w_{i}^{{fus}}$. Subsequently, we introduce a forward gating network $Z_{i}=\{z_{i,r}\}$ into the bypass of the forward propagation, where each gate is controlled by a sigmoid function and corresponds to one layer $r$ of the network. The forward propagation of each layer can then be formulated as
$w_{i,r}^{eff}=w_{i,r}^{{fus}}+z_{i,r}\Delta w_{i,r},$ (28)
where $\Delta w_{i,r}$ denotes the parameter updates during training. When $z_{i,r}=0$, the parameters of this layer are frozen and remain identical to those obtained after Phase 1. Conversely, when $z_{i,r}=1$, the layer is unfrozen and participates in training.
During model optimization, we adopt a two-stage alternating update scheme consisting of an inner and an outer loop. In the inner loop, the gating network $Z_{i}$ is frozen while only the fused model $w_{i}^{{fus}}$ is updated using a small amount of data. This stage requires only a few update steps, allowing the gradients to be accumulated and stored within the gating network. Subsequently, in the outer loop, the model is fixed and the gating network is updated. The optimization objective at this stage is defined as
$J=\mathcal{L}_{acc}(w_{i}^{{eff}})-\beta\mathcal{L}_{rob}(w_{i}^{{eff}})+\lambda_{3}{\left[\sum_{r}z_{i,r}-B\right]}_{+},$ (29)
where $B$ denotes the total training budget, and $\cdot$$_{+}$ represents a penalty term that is activated only when the budget is exceeded and remains zero otherwise. This objective encourages the gating network to identify a subset of layers that reduce robustness loss but increase accuracy loss, selectively opening them for training while keeping the total number of active layers within the budget. After each outer-loop update, the model parameters are reset to $w_{i}^{fus}$ and further trained for multiple iterations.
Through the alternating optimization, the gating network gradually converges to the target $J$. In the final outer-loop iteration, a top-$B$ projection is applied to retain the $B$ largest gates, forming the selected layer set $\mathbf{R}=\{r_{1},r_{2},\dots,r_{B}\}$, where the corresponding layers are opened for training while the others remain frozen. The model is then retrained on clean data for final fine-tuning.
III-E The Workflow of Lorica
Algorithm 1 Lorica’s Adversarial Training Process
@@ III-E The Workflow of Lorica
1:Input: Aggregation epoch $T_{1}$, local training epoch $T_{2}$, local benign data $\mathcal{D}_{i}$, global model $w_{g}(w_{g}^{L},w_{g}^{P},w_{g}^{C})$, and local model $w_{i}(w_{i}^{L},w_{i}^{P},w_{i}^{C})$, where $i\in\{1,2,\dots,N\}$.
@@
3:Initialize: Server’s model $w_{g}$ and clients’ models $w_{i}$.
4:for 11 to $T_{1}$ do
5:  for client $i\in\{1,2,\dots,N\}$ parallel do
6:   Downloads the global LoRA-FA as $w_{i}^{L}$ and expert model $w_{e,i}^{L}$. Combine $w_{i}^{L}$, $w_{i}^{P}$ and $w_{i}^{C}$ to form $w_{i}$;
7:   for 1 to $T_{2}$ do
8:     for $x\in\mathcal{D}_{i}$ do
9:      $x_{adv}=\text{PGD}(w_{i},x)$;
10:      Train and calculate loss $\mathcal{L}$ with eq. 17;
11:      $w_{i}^{L},w_{i}^{C}\leftarrow\text{SGD}(w_{i},\mathcal{L})$;
12:     end for
@@
14:   Each client upload LoRA-FA model $w_{i}^{L}$ to server;
15:  end for
16:  Server performs Byzantine-aware aggregate LoRA-FA model with eq. 9;
17:end for
18:for client $i\in\{1,2,\dots,N\}$ parallel do
19:  Fuse $w_{i}^{L}$, $w_{i}^{P}$, and $w_{i}^{C}$ to form the model $w_{i}^{fus}$, initialize forward-gating network $Z_{i}$;
20:  for 1 to $T_{3}$ do
21:    for 1 to $T_{4}$ do
22:     Freeze $Z_{i}$ and train the model by eq. 28;
23:    end for
24:    Freeze the model and optimize $Z_{i}$ by eq. 29;
25:  end for
26:  Select top-$B$ largest gate as $\mathbf{R}=\{r_{1},r_{2},\dots,r_{B}\}$;
27:  Freeze layers except $\mathbf{P}$ and train model $w_{i}^{fus}$ by $\mathcal{D}_{i}$;
28:end for
algorithm 1 outlines the Lorica defense algorithm. Each client $i$ holds its private benign dataset $\mathcal{D}_{i}$. A global model $w_{g}$ and a local model $w_{i}$, both utilizing LoRA-FA for fine-tuning, are constructed and deployed.
In Phase 1, personalized adversarial fine-tuning is conducted as follows. The server initializes the global model $w_{g}$, and clients download the LoRA-FA $w_{g}^{L}$ and expert model $w_{e,i}^{L}$ to build their local models $w_{i}$. Each client generates adversarial examples from its local dataset $\mathcal{D}_{i}$ using the PGD method and trains locally for $T_{2}$ epochs with eq. 17, updating only its LoRA-FA $w_{i}^{L}$ and classifier $w_{i}^{C}$. After local training, clients upload $w_{i}^{L}$ to the server, which aggregates them via eq. 9 to update the global LoRA-FA $w_{g}^{L}$ and expert model $w_{e,i}^{L}$. This process repeats for $T_{1}$ rounds.
After completing Phase 1, each client possesses the shared global LoRA-FA model $w_{i}^{L}$, the frozen pre-trained backbone $w_{i}^{P}$, and a personalized classifier $w_{i}^{C}$. Phase 2 involves forward-gating selection fine-tuning. First, the model is fused and augmented with a bypass gating network, yielding $w_{i}^{\mathrm{fus}}$ and $Z_{i}$. Then, the gating network is iteratively optimized according to eq. 28 and eq. 29, enabling layer-wise selection. Finally, the top-$B$ largest gates are selected as $\mathbf{P}$. All other layers are frozen. The fused model $w_{i}^{\mathrm{fus}}$ is further trained.
Through this workflow, each client obtains a personalized model suited to its data distribution, achieving high benign accuracy while maintaining strong adversarial robustness.
TABLE II: Performance comparison of Lorica and baseline under different datasets and attack algorithms (ViT-B/16)
@@
Sylva 26.95 27.63 26.23 28.95 28.70
Lorica(Ours) 27.93 28.45 27.77 29.43 29.15
IV Experiments
We evaluate Lorica across diverse distributed settings and attack scenarios, comparing it with baselines to assess personalized defense performance and adaptability. RQ1: How does Lorica perform against different attack algorithms in terms of robustness and accuracy? RQ2: How does Lorica handle varying heterogeneous data distributions across diverse clients? RQ3: How does Lorica achieve efficiency in communication and computation in distributed environments? RQ4: How effective is Lorica in defending against Byzantine attacks during adversarial training? RQ5: How do Lorica’s hyperparameters influence its overall robustness and accuracy performance?
IV-A Experimental Setup
@@
Datasets and models. This study leverages datasets widely recognized in adversarial training, including CIFAR-10 [15], STL-10 [5], and GTSRB [34], as introduced in section II. To better align with tasks involving large models, the more complex CIFAR-100 dataset is also incorporated. For model selection, we utilize large vision models based on transformers, specifically ViTs [9] pre-trained on the ImageNet dataset [33]. These models are categorized into three scales—ViT-T/16, ViT-B/16, and ViT-L/16.
@@
Figure 7: Comparison of adversarial training performance under different heterogeneity levels (ViT-B/16)
@@
Attacks. We simulate both white-box and gray-box attacks using recent and widely adopted image attack algorithms. For white-box attacks, we employ the classic FGSM algorithm [10] and its variants, alongside the more robust PGD algorithm [25]. To explore subtle perturbations, we include the SparseFool attack [27], which modifies only a few pixels. For gray-box attacks, we utilize the PAP algorithm [2], which generates universal perturbations tailored for pre-trained models and remains effective against fine-tuned models without requiring knowledge of downstream tasks. We consider Byzantine attackers with partial client visibility rather than global knowledge, including label-flipping attacks [32] at the data level and model poisoning using the MPAF [3].
@@
Baselines. We use several standard defense algorithms for adversarial training: PGD-AT [25], which employs the PGD algorithm to generate adversarial samples and enhance model robustness; TRADES [42], which optimizes the loss to balance accuracy on benign data and robustness against adversarial attacks; and Gen-AF [48], a fine-tuning approach for adversarial training on pre-trained models, maintaining robustness while improving accuracy. In the distributed scenario, we employ FedAvg [26] and FedProx [17], with FedProx addressing client data heterogeneity. We integrate them with various defense algorithms, establishing them as baseline approaches for comparison. Additionally, we incorporate DBFAT [43], designed specifically for non-IID settings, into our baseline approaches. We also include two personalized federated learning algorithms for comparison. [1] proposes addressing Byzantine attacks in distributed training by customizing the loss function for each client, which we refer to as Per-Adv. [37] introduces a method that computes LoRA-based trust weights among clients during large model fine-tuning to achieve personalized updates, which we denote as Per-LoRA. It is worth mentioning that we incorporate the TRADES loss into both personalized algorithms, thereby improving their alignment with the threat model and experimental tasks defined in this study. For Byzantine robustness, we further compare against widely adopted aggregation rules, including Median and Trimmed Mean [41]. In addition, as expected, we also compare our method with our previous work Sylva [28].
@@
Implementation. For data heterogeneity, client data is partitioned using a Dirichlet distribution with 15 clients and a Dirichlet parameter of 10. In Phase 1, $\gamma$ and $\epsilon$ are set to 0.9, $\lambda_{1}$ to 20, $\lambda_{2}$ to 0.001, and $k$, the number of clients for weight aggregation, is set to 5. We set the malicious client ratio $\rho$ to 5–15% only when evaluating Byzantine attacks. In Phase 2, $\beta$ is set to 5, $B$ to 300, and $p$, the proportion of layers selected for training, to 3% of the total layers. The learning rate is 0.005. FAT uses 5 local epochs, with ViT-T performing 30 cloud aggregation rounds, while ViT-B and ViT-L performing 50. Experiments are conducted on 4 NVIDIA A100 GPUs using the Ray framework for multi-process client simulation. To evaluate Lorica’s efficiency in real-world settings, we conduct experiments on five representative devices, including three consumer GPUs (RTX 4090, 3090, and 2080Ti) and two edge platforms for autonomous driving: the Apollo D-KIT Advanced (RTX 3060) and Jetson AGX Orin.
IV-B Performance of Adversarial Training
@@
We evaluate on four datasets, reporting average performance across all clients (table II). In heterogeneous settings, non-IID-aware baselines like FedProx variants and DBFAT show relatively better performance. Lorica consistently outperforms across diverse datasets, excelling on both benign and adversarial samples. On datasets like CIFAR-10, it shows substantial gains over both traditional federated adversarial training and personalized methods. Specifically, Lorica leads to an accuracy increase of up to 55.0% and a robustness enhancement of 35.3%, demonstrating its superior capability in enhancing both performance and resilience against adversarial attacks. When compared to the baselines of the two personalized algorithms, Lorica continues to deliver significant improvements, achieving an average increase of 50.2% in benign accuracy and 18.7% in robustness. This performance boost can be attributed to Lorica’s greater adaptability, as it effectively balances generalization across tasks with maximizing personalization for individual clients, thereby enhancing both accuracy and robustness. Moreover, benefiting from the optimizations of the Lorica modules, both accuracy and robustness show consistent improvements over Sylva. While the improvements on more complex datasets like CIFAR-100 are comparatively smaller, they remain significant, highlighting Lorica’s adaptability and effectiveness in varied scenarios. Furthermore, we evaluate Lorica on models of varying scales, with detailed results provided in Appendix B. Lorica consistently delivers strong performance across all model scales.
@@
Answer to RQ1: By leveraging its personalized design, Lorica enables each client’s model to effectively adapt to its specific data distribution, achieving superior accuracy and robustness even in highly heterogeneous environments.
TABLE III: Efficiency comparison of Lorica on different edge devices
[…]
ViT-T FAT 1.04 0.35 0.45 0.47 0.49 0.99 0.7 1.3 1.5 1.0 1.2
Per-LoRA 0.89 0.31 0.32 0.33 0.32 0.82 0.9 0.7 1.1 1.1 1.0
Sylva 0.89 0.30 0.32 0.33 0.32 0.82 0.8 0.6 1.2 0.9 0.9
Lorica 0.63 0.28 0.30 0.31 0.31 0.77 0.7 0.5 1.1 0.7 0.8
ViT-B FAT 4.86 0.40 0.47 0.49 0.53 1.13 15.0 13.5 13.8 14.1 13.8
Per-LoRA 3.14 0.34 0.39 0.42 0.45 0.92 1.3 1.4 1.4 1.2 1.5
Sylva 3.14 0.33 0.37 0.41 0.45 0.91 0.8 0.9 0.8 0.7 0.8
Lorica 2.73 0.29 0.34 0.37 0.39 0.80 0.8 0.7 0.8 0.6 0.7
ViT-L FAT 11.43 0.75 0.93 1.78 1.87 2.41 49.9 47.6 49.1 48.7 48.2
Per-LoRA 6.63 0.57 0.88 1.29 1.39 1.71 3.1 2.9 3.2 3.4 3.0
Sylva 6.63 0.55 0.87 1.28 1.39 1.70 1.1 1.0 1.2 1.2 1.1
Lorica 5.37 0.43 0.65 0.92 1.01 1.38 0.8 0.7 0.8 0.9 0.9
IV-C Performance of Different Heterogeneous Distributions